MIDC News
Year of 2010
December 15, 2010
- 09:01 MST: January through September 2010 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 09:18 MST: January through September 2010 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 16:00 MST: Ethernet cable to SRRL radiometer tower has been repaired. -Afshin Andreas
- 08:36 MST: SRRL BMS upwelling PIR singal cable was repaired at 16:12 MST yesterday. Data collection to radiometer tower is running temporarily on an rs-232 cable above ground, until ethernet cable repair is complete. -Afshin Andreas
- 09:36 MST: SRRL BMS Data collection is down as of 09:03 MST today. Construction has broken an ethernet cable to the radiometer tower. Also a signal cable to the upwelling PIR has been damaged.
-Afshin Andreas
- 14:32 MST: The SRRL BMS PV Tower was dropped between 11:00 and 14:00 MST today, so that the wind speed sensor at 22 feet could be replaced. The old sensor had a broken cup. -Afshin Andreas
- 10:37 MST: The SRRL BMS Global SP-110 was swapped out with a new sensor today. Logger was updated with new calibration factor at 10:35 MST. -Afshin Andreas
- 10:43 MST: Atmospheric Electric Field was added to the SRRL BMS Graphical Display at 9:30 MST today. -Afshin Andreas
- 13:51 MST: The SRRL BMS instantaneous data stream for the radiometer tower was changed at 13:40 MST. Wind Direction at 6' was replaced with Atmospheric Electric Field. Note: This data stream is for internal use only. -Afshin Andreas
- 07:32 MST: Atmospheric Electric Field was added to the SRRL BMS data stream on September 27. Data value will be added to the graphical display in the future. -Afshin Andreas
- 10:06 MST: Data from the Solar Resource & Meteorological Assessment Project (SOLRMAP): Southwest Solar Technologies SolarCAT - Rotating Shadowband Radiometer (RSR) is now available on MIDC. Data collection started September 27, 2010. -Afshin Andreas
- 08:52 MST: The SRRL BMS HMP-45C Temp/RH probe (data lab) was calibrated and returned to service at 08:44 MST today. -Afshin Andreas
- 16:25 MST: Daily SRRL YES TSI-880 sky animations are now available for both the raw images and the cloud analyzed images. -Afshin Andreas
- 09:27 MST: The following SRRL BMS radiometer tower sensors were calibrated against a LICOR LI-1800 spectroradiometer on 9/9/2010: Global Quantum LI-190, Global Photometric LI-210, and Upwelling Quantum LI-190. Also the Global CM3 was calibrated using BORCAL 2010-03 reference data and the Upwelling CM3 was calibrated against the BMS Upwelling PSP. New factors for all these instruments were entered into the logger @ 9:18 MST today. -Afshin Andreas
- 13:33 MST: The SRRL BMS Upwelling Quantum sensor was facing the sky from 11:25 to 13:11 MST, so that the instrument could be calibrated. Data should be ignored during that time. -Afshin Andreas
- 14:50 MST: The following SRRL BMS radiometer tower instruments were swapped out with instruments that were calibrated in BORCAL 2010-01 (PSPs) and 2010-03 (LI-200s) at approx 13:42 to 14:12 MST today: Global 90-North/South/East/West PSPs, Global 90-East/West LI-200s, and Global 40-South LI-200. Cal factors were updated in the logger @ 14:03 MST today -Afshin Andreas
- 12:43 MST: The SRRL PGS-100 spectroradiometer was calibrated on August 20, 2010 outdoors against the NIST traceable LI-1800 PRS-158. The new calibration file has been applied for data starting on September 3, 2010. This calibration also included a wavelength calibration. -Afshin Andreas
- 13:27 MST: The SRRL BMS CR23x data logger was calibrated at 12:24 to 13:16 MST today. The following instruments may read bogus during that time: Diffuse Shadowband PSP, Direct LI-201, Global SP-Lite, Global SP-110, Global TSP-1. -Afshin Andreas
- 17:00 MST: The SRRL BMS HMP-45C Temp/RH probe (datalab sensor) was removed at 10:16 MST for calibration. -Afshin Andreas
- 07:31 MST: The SRRL BMS TSP-700 has had a noise problem since August 11, instrument will be troubleshot later this month. -Afshin Andreas
- 16:06 MST: Redundant RAID level 1 hard drive was replaced on internal MIDC server today. -Afshin Andreas
- 10:38 MST: MIDC maintenance records, data quality statements (DQS), and calibration histories are down again due to problems that starting occuring after public server relocations at NREL. -Afshin Andreas
- 11:48 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service today @ 9:42 MST, calibrated against NIST F620, and returned to service @ 11:28 MST. -Afshin Andreas
- 13:56 MST: Removed faulty bubble level from SRRL BMS SP-Lite and re-leveled instrument. -Afshin Andreas
- 09:31 MST: The CS107 temperature probe on the SRRL RSR2 was replaced with WXT520 today @ 09:07 MST. The WXT520 sensor includes: temperature, relative humidity, wind speed & direction, pressure, and rail/hail. The WXT520 is for testing only and may be removed in the future. -Afshin Andreas
- 11:14 MST: Atmospheric Electric Field was added to the NWTC M2 Graphical Display at 11:06 MST today. -Afshin Andreas
- 09:21 MST: The SRRL BMS Global CM22 was calibrated in BORCAL 2010-03. The new cal factor was updated in the logger program today @ 9:20 MST. -Afshin Andreas
- 10:47 MST: The original SRRL BMS CH1 S/N 010256 was returned to service today. Instrument was resealed and calibrated in BORCAL 2010-02. Also the BMS SPN1 was replaced with a new instrument S/N A360 using the factory calibration. Data logger was updated at 9:47 MST today for both instruments. -Afshin Andreas
- 11:10 MST: MIDC maintenance records, data quality statements (DQS), and calibration histories are now operational. -Afshin Andreas
- 10:06 MST: MIDC maintenance records, data quality statements (DQS), and calibration histories have been down since July 22 due to public server relocations at NREL. Estimated downtime is unknown. -Afshin Andreas
- 13:29 MST: Atmospheric Electric Field was added to the NWTC M2 data stream at 13:05 MST today. Data value will be added to the graphical display in the near future. -Afshin Andreas
- 11:57 MST: Data from the Solar Resource & Meteorological Assessment Project (SOLRMAP): Utah Geological Survey - State Energy Program - Rotating Shadowband Radiometer (RSR) for both Cedar City and Milford are now available on MIDC. Data collection started July 13, 2010 and July 14, 2010, respectively. -Afshin Andreas
- 13:35 MST: Due to several tripped GFCI outlets at SRRL the following instruments were without power from July 4 @ 18:10 to July 6 @ 7:30 MST: Global CUVA1, CUVB1, MS210W, 501A and Direct CUVA2, CUVB2, PGS100. -Afshin Andreas
- 09:36 MST: The following SRRL BMS deck instruments were swapped out with instruments that were calibrated in BORCAL 2010-01 at approx 8:36 to 8:40 MST today: Global RG780 PSP, Zebra PSP, Direct RG780 NIP, Shadowband PSP, Diffuse PSP, and Diffuse 8-48. Cal factors were updated in the logger @ 8:40 MST today. -Afshin Andreas
- 14:24 MST: The following SRRL BMS radiometer tower instruments were swapped out with instruments that were calibrated in BORCAL 2010-01 at approx 12:46 to 12:56 MST today: Global 40-South PSP, Upwelling PSP, Global Normal CM-21, Global 90-North LI-200, Global 90-South LI-200, and Upwelling LI-200. Cal factors were updated in the logger @ 12:50 MST today. -Afshin Andreas
- 11:13 MST: The SRRL BMS Global CM22 and Diffuse CM22 instruments were swapped today between 10:35 and 10:40 MST, so that the diffuse instrument can be calibrated in the next BORCAL. Logger program was updated at 10:41 MST. -Afshin Andreas
- 12:23 MST: A lightning strike at South Park Mountain Data damaged the wind speed & direction sensor as well as two temperature sensors on June 13. The wind speed & direction have been permanently removed from site. Temperature sensors will be replaced at a later date. Data stream format was updated today, all historical data since June 14 was converted to the new format. -Afshin Andreas
- 11:06 MST: The following SRRL BMS instruments were calibrated in BORCAL 2010-01: Global Horizontals: CM22, CM6b, PSP (ventilated), PSP, TSP-700, TSP-1, SP-Lite, LI-200, SP-110, SPN1; Direct Normals: CH1, NIP #2, NIP #1, LI-201; Diffuse SPN1. Logger program was updated at 10:32 MST today with new cal factors. Responsivity Coefficients for Global PSPs were also updated at 11:01 MST today. -Afshin Andreas
- 11:06 MST: Due to the water leak in the SRRL BMS CH1, the instrument was swapped with a loaner instrument S/N 080033 at 10:08 MST today. Calibration factor was updated in logger program at 10:32 MST. -Afshin Andreas
- 14:14 MST: The file storage method for the SRRL TSI images has been upgraded. Also a new link is available when viewing an image that allows the user to download an entire day's worth of images in one ZIP file. -Afshin Andreas
- 12:36 MST: The loaner PGS-100 was installed again at SRRL today at 12:20 MST. The loaner is now owned by NREL and the old unit will be returned to manufacture. -Afshin Andreas
- 10:48 MST: The SRRL PGS-100 Direct Normal Spectroradiometer shutter is sticking closed again (as expected). -Afshin Andreas
- 08:27 MST: There is a large night time offset (up to -40 W/m^2) on the BMS Global 90-West PSP which started April 23 @ 16:20 MST. Problem will be investigated in the near future. -Afshin Andreas
- 10:53 MST: The loaner PGS-100 was replaced with the NREL owned unit at 10:50 MST today. Factory found no problem with the shutter. -Afshin Andreas
- 08:08 MST: There is a problem with the SRRL BMS Global MS210W which started last night. The signal is negative, problem will be investigated in the near future. -Afshin Andreas
- 07:09 MST: Data from the Solar Resource & Meteorological Assessment Project (SOLRMAP): Loyola Marymount University - Rotating Shadowband Radiometer (RSR) is now available on MIDC. Data collection started April 6, 2010. -Afshin Andreas
- 14:42 MST: Data from the Solar Resource & Meteorological Assessment Project (SOLRMAP): Kalaeloa Oahu, Hawaii - Rotating Shadowband Radiometer (RSR) is now available on MIDC. Data collection started March 16, 2010. -Afshin Andreas
- 16:06 MST: The SRRL PGS-100 factory loaner was calibrated on Feb 26, 2010 outdoors against the NIST traceable LI-1800 PRS-158. The new calibration file has been applied to all historical data since the loaner unit was installed on Jan 29, 2010. -Afshin Andreas
- 10:04 MST: The OTF Global 40-South MS-700 Spectroradiometer was calibrated against NIST F620 on February 26 and returned to service on March 10 @ 10:08 MST. -Afshin Andreas
- 14:58 MST: The SRRL Global 40-South LI-1800 Spectroradiometer was removed from service today @ 11:26 MST, calibrated against NIST F620, and returned to service @ 13:04 MST. -Afshin Andreas
- 10:13 MST: The OTF MS-700 Spectroradiometer was removed from service on Feb 17 @ 12:15 MST for calibration. -Afshin Andreas
- 14:49 MST: A factory loaner PGS-100 Spectroradiometer was installed today @ 14:40 MST. This loaner is a functional check and is for troublshooting purposes only. Data is not calibrated and should not be used. -Afshin Andreas
- 15:59 MST: The SRRL PGS-100 Direct Normal Spectroradiometer was removed from service today @ 15:50 MST. Unit will be sent back to factory again for repair. -Afshin Andreas
- 11:46 MST: December 2009 Monthly Data Sets & Reports/Plots for SRRL BMS have been generated. -Afshin Andreas
- 11:12 MST: December 2009 Monthly Data Sets & Reports for ECSU and BSC have been generated. -Afshin Andreas
- 08:29 MST: The SRRL BMS HMP-45C Temp/RH probe (PV tower) was swapped out with a calibrated sensor @ 16:05 through 16:07 MST yesterday. -Afshin Andreas
- 15:46 MST: The SRRL PGS-100 Direct Normal Spectroradiometer shutter is sticking closed again (causing no signal). First intermittent occurance was on November 18, 2009. -Afshin Andreas
News Archives:
[1997] [1998] [1999] [2000] [2001] [2002] [2003] [2004] [2005] [2006] [2007] [2008] [2009] [2010] [2011] [2012] [2013] [2014] [2015] [2016] [2017] [2018] [2019] [2020] [2021] [2022] [2023] [2024] [2025]